# What to align in multimodal contrastive learning?

### v1

@@ Abstract
Humans perceive the world through multisensory integration, blending the information of different modalities to adapt their behavior. Contrastive learning offers an appealing solution for multimodal self-supervised learning. Indeed, by considering each modality as a different view of the same entity, it learns to align features of different modalities in a shared representation space. However, this approach is intrinsically limited as it only learns shared or redundant information between modalities, while multimodal interactions can arise in other ways. In this work, we introduce CoMM, a Contrastive Multimodal learning strategy that enables the communication between modalities in a single multimodal space. Instead of imposing cross- or intra- modality constraints, we propose to align multimodal representations by maximizing the mutual information between augmented versions of these multimodal features. Our theoretical analysis shows that shared, synergistic and unique terms of information naturally emerge from this formulation, allowing us to estimate multimodal interactions beyond redundancy. We test CoMM both in a controlled and in a series of real-world settings: in the former, we demonstrate that CoMM effectively captures redundant, unique and synergistic information between modalities. In the latter, CoMM learns complex multimodal interactions and achieves state-of-the-art results on the six multimodal benchmarks.
@@ 1 Introduction
Multisensory or multimodal learning (Baltrušaitis et al., 2018) involves extracting and processing information from multiple sources (or modalities, e.g. text, audio, images, tabular data, etc.) to perform a task. The whole human experience is inherently multimodal: we simultaneously see, hear, smell, taste and feel, and these different sensory signals are combined to give us the necessary information to explore our environment. Indeed, many of the simplest tasks we tackle in our daily lives are multimodal. For example, the way we perceive the flavor of our food or drinks does not depend solely on our taste, but also on what we see (Morrot et al., 2001) or even what we hear (Woods et al., 2011) while we eat. McGurk & MacDonald (1976) have also demonstrated that visual stimuli interact with audio signals to perform human speech recognition.
@@
We propose a Contrastive MultiModal self-supervised pre-training method (CoMM) that enables the communication of modalities in a single multimodal space. Unlike previous contrastive multimodal methods that impose cross-modal constraints to align unimodal representations, we propose to leverage a simple multimodal architecture to fuse multimodal inputs into a common representation and then align the multimodal features by maximizing the mutual information between augmented versions of these features (see Fig. 1a). CoMM enables to model multimodal interactions –including redundancy, uniqueness and synergy– in the context of multimodal representation learning for the first time111to the best of our knowledge , as these terms naturally emerge from our contrastive multimodal formulation.
@@
Based on Partial Information Decomposition (Williams & Beer, 2010; Bertschinger et al., 2014), we built a strong theoretical basis for CoMM to learn multimodal interactions in a self-supervised way. Empirically, we show that CoMM effectively captures redundant, unique and synergistic information between modalities in a controlled environment, where the type of interaction is known. In a series of real-world datasets –from different domains (healthcare, robotics, etc.) and including diverse data types (image, text, audio, etc.)–, CoMM achieves state-of-the-art results on seven multimodal tasks with two or three modalities. Thus, CoMM shows to be a versatile framework capable of handling any number of modalities, various data types, and different domains.
@@ 2 Quantifying multimodal interactions
Problem setup.
Let $X_{1},X_{2},\ldots,X_{n}$ be random variables representing $n$ different data modalities (e.g. images, text, audio, tabular data, etc.), and a given task $Y$. Our goal is to learn a latent variable $Z=f(X)$ that is a good representation of $X=(X_{1},\ldots,X_{n})$ for $Y$.
For our theoretical analysis, we set $n=2$, as multimodal interactions have not been characterized yet by PID for larger $n$. All proofs can be found in Appendix G.
@@
Existing methods using contrastive objectives to learn multimodal representations (Jia et al., 2021; Radford et al., 2021; Tian et al., 2020) impose cross-modal constraints by maximizing an estimator of $I(X_{1};X_{2})$ as an approximation of $I(X_{1},X_{2};Y)$. However, this strategy is limited by the multiview redundancy assumption.
@@
What happens when other sources of multimodal information intervene? FactorCL (Liang et al., 2023b) is a good initial step to integrate uniqueness and redundancy into multimodal contrastive learning by applying multimodal augmentations. However, it heavily relies on the assumption that optimal multimodal augmentations can be obtained by applying a two-step process based on conditional augmentations. We argue that this hypothesis is unrealistic, as the first unique augmentation is strongly related to task-relevant information of different modalities. For example, if the text caption is “a yellow flower”, then color jittering should not be applied to an image depicting a flower. Besides, the factorized formulation of FactorCL is impractical as it is prone to cumulative errors. Finally, this method does not consider the synergy between modalities. In contrast, we propose a model that relies solely in the main hypothesis of contrastive learning, extended to the multimodal case, without relying on strong assumptions between multimodal relationships nor conditional augmentations.
Assumption 1 (Label-preserving multimodal augmentations)
@@
Even if 1 might seem strong at first glance, it actually makes sense in the context of multimodal representation learning. Indeed, coming back to the example of the flower image with “a yellow flower” as a caption, applying color jittering to the image would allow the model to focus on other interactions (uniqueness or synergy) rather than color redundancy or even to refocus on other features (like the flower shape).
@@ 3.2 Multimodal architecture
Transformer block. The goal of this module is to perfom the fusion of the modality-specific embeddings through the multihead self-attention layers in the transformer block, obtaining the final multimodal embedding $Z$.
@@ 3.3 Training
These terms are encoded by the network to obtain $(n+2)$ embeddings, namely: $Z^{\prime},Z^{\prime\prime}$, and $\{Z_{i}\}_{i=1}^{n}$. $(2n+1)$ mutual information terms are then optimized through backpropagation: $I(Z^{\prime},Z^{\prime\prime})$ to maximize $I(Z,Z^{\prime})$ in Eq. 3 and $I(Z_{i},Z^{\prime})$ and $I(Z_{i},Z^{\prime\prime})$ to better approximate $R+U_{i}$ in Eq. 5 for $i\in\{1,...,n\}$.
@@ 4 Experiments
Evaluation. Given a pre-trained model $f$, we perform evaluation through linear probing, i.e., we train a linear layer $g_{W}(x)=Wf(x)$ ($f$ fixed) to minimize the classification or regression loss (depending on the task); and fine-tuning, further training $f$ in a supervised way.
4.1 Controlled experiments on bimodal Trifeature dataset
Figure 4: Linear probing accuracy of redundancy (shape), uniqueness (texture) and synergy (color and texture) on bimodal Trifeature dataset. CoMM is the only model capturing all three task-related interactions between modalities [^4] We ran the experiments with the official implementation of FactorCL (Liang et al., 2023b) but performances were close to random chance so we did not integrate them to this analysis..
Does CoMM learn redundant, synergistic and unique information for a given task? We design a synthetic bimodal dataset based on Trifeature (Hermann & Lampinen, 2020) to answer. Briefly, we generate a Trifeature dataset (as first modality) containing images of one of ten shapes, for one of ten textures and one of ten colors (1000 combinations in total). We augment each image three times using rotations and translations. Then, we pair each image with a second one (as second modality) from the same Trifeature dataset, allowing us to control the shared, unique and synergistic attributes between the two modalities. As Hermann & Lampinen (2020), we use AlexNet (Krizhevsky et al., 2012) as modality-specific encoder for both modalities in all experiments with an embedding dimension $d=512$.
@@ 4.1 Controlled experiments on bimodal Trifeature dataset
In Fig. 4, we show that cross-modality constraints with InfoNCE (Radford et al., 2021) (“Cross” model) allow to perfectly capture redundant information but completely remove unique and synergistic information, as predicted in Lemma 1. Self-supervised constraints on each encoder (“Cross+Self” (Yuan et al., 2021)) capture accurately unique information but fail at preserving synergy. CoMM is the only model learning all three interactions related to the task.
4.2 Experiments on real-world datasets with 2 modalities
@@
Multimodal IMDb (MM-IMDb) (Arevalo et al., 2017) is a dataset for movie genre prediction. We consider two modalities: images and text (the movie poster and its plot, respectively). Since each movie can be classified into one or more genres, it is a multi-label classification task, with 23 categories. MM-IMDb provides a suitable example of life-like multimodal task as the genre prediction cannot be performed accurately from the movie poster or the movie plot alone, while results significantly improve by considering both (Arevalo et al., 2017), suggesting that synergistic interactions are needed to fulfill this task. We compare CoMM’s performance on MM-IMDb against important baselines under two different settings. First, in the self-supervised setting, we consider CLIP (Radford et al., 2021), representing “Cross” methods, SLIP (Mu et al., 2022) representing “Cross+Self” methods, SimCLR (Chen et al., 2020a) for unimodal self-supervised methods, BLIP-2 (Li et al., 2023) as a recent powerful vision and language model, as baselines. All models were trained on unlabeled MM-IMDb. For CLIP, we also report results with the publicly released weights without further training. CoMM is initialized with pre-trained weights from CLIP and BLIP-2. Second, in the fine-tuning setting, we compare CoMM fine-tuned with state-of-the-art fully supervised baselines. We also include results for LLaVA-NeXT (Li et al., 2024) representing new vision-and-language generative models, however these scores are not fully comparable to ours, since the model might have seen IMDb database during training. See Appendix B for implementation details.
@@
Table 2 shows that CoMM outperforms all models in both settings. In the self-supervised setting, CoMM has a margin of $7.5\%$ and $5.8\%$ of macro and weighted F1-scores, respectively, with the second-best method. It is interesting to observe that CLIP performs better with the publicly released weights (probably because of the large-scale and diversity of the data it was trained on), than with further training on MM-IMDb. This result suggests that the genre movie prediction does not benefit from learning redundant information. Including uniqueness from the image modality allows for some improvement (SLIP). In the fine-tuning setting, CoMM fine-tuned outperforms existing fully supervised baselines even if MFAS has been designed to search for the best fusion strategy.
4.3 Experiments on real-world datasets with 3 modalities
@@
Table 3: Linear evaluation top-1 accuracy (%) on Vision&Touch (Lee et al., 2020) in the robotics domain and UR-FUNNY.
@@
In Table 3, we compare CoMM trained on the three modalities in a self-supervised way against CMC (Tian et al., 2020). We also compare with bi-modal models: CoMM, Cross and Cross+Self trained on image and proprioception data for Vision&Touch and image and text data for UR-FUNNY.
First, we observe a consistent improvement ($+6.1\%$ on V&T, $+1.5\%$ on UR-FUNNY) when adding a third modality with CoMM (compared to only using two), which demonstrates its versatility. Second, we improve the state-of-the-art for SSL methods on datasets with more than two modalities ($+0.1\%$ and $+5.4\%$ on V&T and UR-FUNNY respectively).
@@ 5 Ablation studies
Loss function. First, we check our claim that optimizing both $\mathcal{L}$ and $\sum_{i=1}^{n}\mathcal{L}_{i}$ –as in Eq. 7– is required to accurately capture uniqueness, synergy, and redundancy. In  Fig. 5, we show that maximizing $\sum_{i=1}^{n}\mathcal{L}_{i}$ improves redundancy and uniqueness, as guaranteed by our Lemma 3, but fails for synergy. Conversely, maximizing $\mathcal{L}$ allows one to learn all information terms but very slowly. In particular, for synergy, we argue it is because the model has to learn modality-specific features first (phase 1) before learning their interactions (phase 2). The former is learned through $I(Z_{i},Z^{\prime})+I(Z_{i},Z^{\prime\prime})$ while the latter is captured with $I(Z^{\prime},Z^{\prime\prime})$. Hence, $\mathcal{L}_{\text{CoMM}}$ speeds up phase 1 and can learn synergy more efficiently in phase 2.
@@
CoMM $\mathbf{99.92}_{\pm 0.03}$ $\mathbf{87.83}_{\pm 1.55}$ $\mathbf{71.87}_{\pm 2.06}$ $\mathbf{86.83}$
@@ 6 Related work
Multimodal learning refers to methods that connect and integrate information from multiple sources of data (Baltrušaitis et al., 2018; Akkus et al., 2023). Early works focused on training separate encoders and studied different fusion mechanisms to blend the information from different inputs (Zeng et al., 2007; Pérez-Rúa et al., 2019). With the development of transformers and ViTs, the focus has shifted towards training a unique architecture and designing specific tasks and loss functions to integrate multimodal interactions (Xu et al., 2023; Lu et al., 2019; Sun et al., 2019; Chen et al., 2020b; Lu et al., 2020). Today multimodal learning includes different research lines, from generative multimodal learning (Suzuki & Matsuo, 2022; Ramesh et al., 2021; Saharia et al., 2022; Wang et al., 2022b; Alayrac et al., 2022) to multimodal representation learning (Zong et al., 2023; Tian et al., 2020). CoMM belongs to the latter category.
@@
Self-supervised multimodal representation learning. Self-supervised learning aims to learn general representations through supervisory signals from the data itself (known as the pretext task) (Balestriero et al., 2023). In the multimodal context (Zong et al., 2023), self-supervised methods can be grouped according to their pretext task. Clustering-based methods (Alwassel et al., 2020; Asano et al., 2020), where cluster assignments are used as pseudo-labels to train the model and different modalities can be used as supervisory signals to each other; masking modeling (Mizrahi et al., 2023; Bachmann et al., 2023; Lu et al., 2023) methods that reconstruct pieces of information that have been masked from input data. In the multimodal case, masked modelling is used in a cross-modal way, by predicting missing information conditioned in other modalities. Contrastive methods (Radford et al., 2021; Jia et al., 2021) in the multimodal context have been mostly used to matched data from different modalities to obtain aligned representations. In this work, we take a different perspective for contrastive multimodal methods for learning multimodal representations and explore this family of approaches in the following paragraph.
@@
Multimodal contrastive learning. Current methods in contrastive learning are inspired by the idea of multiview learning (Li et al., 2018) and have shown remarkable results in representation learning in unimodal tasks (Chen et al., 2020a; He et al., 2020; Tian et al., 2020; Oord et al., 2018). The natural extension of these approaches to multimodal settings is to optimize a cross-modal contrastive loss (Radford et al., 2021; Alayrac et al., 2020; Jia et al., 2021). Other works have gone even further by introducing cross-modal and intra-modal contrastive objectives (Mu et al., 2022; Jain et al., 2021), or by adding other intra-modality regularization terms (Wang et al., 2022a; Kim et al., 2022). However, these approaches are designed to learn redundant information, neglecting the contributions of uniqueness or synergy . Recently, FactorCL (Liang et al., 2023b) has proposed a solution to model shared and unique task-relevant information explicitly. Yet, the method relies heavily on assumptions that are hard to meet in practice; it proposes a factorized approximation of multimodal interactions that is prone to cumulative errors and does not model synergistic information. Alternatively, we propose CoMM, a contrastive multimodal approach that leverages simple ideas to capture unique, redundant and synergistic interactions.

### v2

@@ Abstract
Humans perceive the world through multisensory integration, blending the information of different modalities to adapt their behavior. Alignment through contrastive learning offers an appealing solution for multimodal self-supervised learning. Indeed, by considering each modality as a different view of the same entity, it learns to align features of different modalities in a shared representation space. However, this approach is intrinsically limited as it only learns shared or redundant information between modalities, while multimodal interactions can arise in other ways. In this work, we introduce CoMM, a Contrastive Multimodal learning strategy that enables the communication between modalities in a single multimodal space. Instead of imposing cross- or intra- modality constraints, we propose to align multimodal representations by maximizing the mutual information between augmented versions of these multimodal features. Our theoretical analysis shows that shared, synergistic and unique terms of information naturally emerge from this formulation, allowing to estimate multimodal interactions beyond redundancy. We test CoMM both in a controlled and in a series of real-world settings: in the former, we demonstrate that CoMM effectively captures redundant, unique and synergistic information between modalities. In the latter, we show that CoMM learns complex multimodal interactions and achieves state-of-the-art results on seven multimodal tasks. Code is available here.
@@ 1 Introduction
Multisensory or multimodal learning (Baltrušaitis et al., 2018) involves extracting and processing information from multiple sources (e.g. text, audio, images, tabular data, etc.). The whole human experience is inherently multimodal: we simultaneously see, hear, smell, taste and feel, and these different sensory signals are combined to give us the necessary information to explore our environment. Many of the simplest tasks we tackle in our daily lives are multimodal:the way we perceive the flavor of our food or drinks does not depend solely on our taste, but also on what we see (Morrot et al., 2001) or what we hear (Woods et al., 2011) while we eat. McGurk & MacDonald (1976) have also shown that visual stimuli interact with audio signals to perform human speech recognition.
@@
To achieve task-agnostic multimodal learning, we propose a Contrastive MultiModal self-supervised pre-training method (CoMM) that enables the communication of modalities in a single multimodal space. Unlike previous contrastive multimodal methods that impose cross-modal constraints to align unimodal representations, we propose to leverage a simple multimodal architecture to fuse multimodal inputs into a common representation and then align the multimodal features by maximizing the mutual information between augmented versions of these features (see Fig. 1a). CoMM enables to model multimodal interactions –including redundancy, uniqueness and synergy– in the context of multimodal representation learning, as these terms naturally emerge from our contrastive multimodal formulation.
@@
Based on Partial Information Decomposition (Williams & Beer, 2010; Bertschinger et al., 2014), we built a strong theoretical basis for CoMM to learn multimodal interactions in a self-supervised way. Empirically, we show that CoMM effectively captures redundant, unique and synergistic information between modalities in a controlled environment, where the type of interaction is known.
Then, in a series of real-world datasets –from different domains (healthcare, robotics, etc.) and including diverse data types (image, text, audio, etc.)–, CoMM achieves state-of-the-art results on seven multimodal tasks with two or three modalities. In all cases, CoMM showed to be a versatile framework capable of handling any number of modalities, various data types, and different domains.
@@ 2 Quantifying multimodal interactions
Problem setup. Let $X_{1},X_{2},\ldots,X_{n}$ be random variables representing $n$ different data modalities (e.g. images, text, audio, tabular data, etc.), and a given task $Y$. Our goal is to learn a latent variable $Z=f(X)$ that is a good representation of $X=(X_{1},\ldots,X_{n})$ for $Y$.
For our theoretical analysis, we set $n=2$, as multimodal interactions have not been characterized yet by PID for larger $n$. In practice, CoMM’s implementation for $n>2$ is straightforward and tested in Section 4.3. All proofs can be found in Appendix G.
@@
Existing methods using contrastive objectives to learn multimodal representations (Jia et al., 2021; Radford et al., 2021; Tian et al., 2020a) impose cross-modal constraints by maximizing an estimator of $I(X_{1};X_{2})$ as an approximation of $I(X_{1},X_{2};Y)$. However, this strategy is limited by the multiview redundancy assumption.
@@
What happens when other sources of multimodal information intervene? FactorCL (Liang et al., 2023b) is a good initial step to integrate uniqueness and redundancy into multimodal contrastive learning by applying multimodal augmentations. However, it heavily relies on the assumption that optimal multimodal augmentations can be obtained by applying a two-step process based on conditional augmentations. We argue that this hypothesis is unrealistic, as the first unique augmentation is strongly related to task-relevant information of different modalities. For example, if the text caption is “a yellow flower”, then color jittering should not be applied to an image depicting a flower. Besides, the factorized formulation of FactorCL is impractical as it is prone to cumulative errors. Finally, this method does not consider the synergy between modalities.
In contrast, we propose a model that relies solely in the main hypothesis of contrastive learning, extended to the multimodal case, without relying on strong assumptions about multimodal relationships nor conditional augmentations.
Assumption 1 (Minimal label-preserving multimodal augmentations)
@@
Even if 1 might seem strong at first glance, it actually makes sense in the context of multimodal representation learning. Indeed, coming back to the example of the flower image with “a yellow flower” as a caption, applying color jittering to the image would allow the model to focus on other interactions (uniqueness or synergy) rather than color redundancy or even to refocus on other features (like the flower shape). This is discussed more in-depth in Section C.4.
@@ 3.2 Multimodal architecture
Transformer block. The goal of this module is to perform the fusion of the modality-specific embeddings through the multihead self-attention layers in the transformer block, obtaining the final multimodal embedding $Z$.
@@ 3.3 Training
These terms are encoded by the network to obtain $(n+2)$ embeddings, namely: $Z^{\prime},Z^{\prime\prime}$, and $\{Z_{i}\}_{i=1}^{n}$. $(2n+1)$ mutual information terms are then optimized through backpropagation: $I(Z^{\prime},Z^{\prime\prime})$ to maximize $I(Z,Z^{\prime})$ in Eq. 3 and both $I(Z_{i},Z^{\prime})$ and $I(Z_{i},Z^{\prime\prime})$ to better approximate $R+U_{i}$ in Eq. 5 for $i\in\{1,...,n\}$.
@@ 4 Experiments
Evaluation. Given a pre-trained model $f$, we perform evaluation through linear probing, i.e., we train a linear layer $g_{W}(x)=Wf(x)$ ($f$ fixed) to minimize the classification or regression loss (depending on the task). We also report results obtained with fine-tuning, i.e. after further training $f$ in a supervised way.
4.1 Controlled experiments on the bimodal Trifeature dataset
Figure 4: Linear probing accuracy of redundancy (shape), uniqueness (texture) and synergy (color and texture) on bimodal Trifeature dataset. CoMM is the only model capturing all three task-related interactions between modalities.
To evaluate whether CoMM learns redundant, synergistic and unique information for a given task, we design a synthetic bimodal dataset based on Trifeature (Hermann & Lampinen, 2020). Briefly, we generate a Trifeature dataset (as first modality) containing images of one of ten shapes, for one of ten textures and one of ten colors (1 000 combinations in total). We augment each image three times using rotations and translations. Then, we pair each image with a second one (as second modality) from the same Trifeature dataset, allowing us to control the shared, unique and synergistic attributes between the two modalities. As Hermann & Lampinen (2020), we use AlexNet (Krizhevsky et al., 2012) as modality-specific encoder for both modalities in all experiments with an embedding dimension $d=512$.
@@ 4.1 Controlled experiments on the bimodal Trifeature dataset
In Fig. 4, we show that cross-modality constraints with InfoNCE (Radford et al., 2021) (“Cross” model) allow to perfectly capture redundant information but completely remove unique and synergistic information, as predicted in Lemma 1. Self-supervised constraints on each encoder (“Cross+Self” (Yuan et al., 2021)) capture accurately unique information but fail at preserving synergy, FactorCL, the method most closely related to our work, also performs poorly on synergy. CoMM is the only model learning all three interactions related to the task.
4.2 Experiments with 2 modalities on real-world datasets
@@
Multimodal IMDb (MM-IMDb) (Arevalo et al., 2017) is a dataset for movie genre prediction. We consider two modalities: images and text (the movie poster and its plot’s description, respectively). Since each movie can be classified into one or more genres, it is a multi-label classification task, with 23 categories. MM-IMDb provides a suitable example of life-like multimodal task as the genre prediction cannot be performed accurately from the movie poster or the movie plot alone, while results significantly improve by considering both (Arevalo et al., 2017), suggesting that synergistic interactions are needed to fulfill this task. We compare CoMM’s performance on MM-IMDb against important baselines under two different settings. First, in the self-supervised setting, we consider CLIP (Radford et al., 2021), representing “Cross” methods, SLIP (Mu et al., 2022) representing “Cross+Self” methods, SimCLR (Chen et al., 2020a) for unimodal self-supervised methods, BLIP-2 (Li et al., 2023) as a recent powerful vision and language model, as baselines. All models were trained on unlabeled MM-IMDb. For CLIP, we also report results with the publicly released weights without further training. CoMM is initialized with pre-trained weights from CLIP and BLIP-2. Second, in the fine-tuning setting, we compare CoMM fine-tuned with state-of-the-art fully supervised baselines. We also include results for LLaVA-NeXT (Li et al., 2024) representing new vision-and-language generative models, however these scores are not fully comparable to ours, since the model might have seen the IMDb database during training. See Appendix B for implementation details.
Table 2 shows that CoMM outperforms all models in both settings. In the self-supervised setting, CoMM has a margin of $7.5\%$ and $5.8\%$ of macro and weighted F1-scores, respectively, with the second-best method. It is interesting to observe that CLIP performs better with the publicly released weights (probably because of the large-scale and diversity of the data it was trained on), than with further training on MM-IMDb. This result suggests that the genre movie prediction does not benefit from learning redundant information. Including uniqueness from the image modality allows for some improvement (SLIP). In the fine-tuning setting, CoMM fine-tuned outperforms existing fully supervised baselines, even if MFAS has been designed to search for the best fusion strategy, and vision-language generative models (LLaVA-NeXT).
4.3 Experiments with 3 modalities on real-world datasets
@@
Table 3: Linear evaluation top-1 accuracy (%) on Vision&Touch and UR-FUNNY.
@@
In Table 3, we compare CoMM trained on the three modalities in a self-supervised way against CMC (Tian et al., 2020a). We also compare with bi-modal models: CoMM, Cross and Cross+Self trained on image and proprioception data for Vision&Touch and image and text data for UR-FUNNY.
First, we observe a consistent improvement ($+6.1\%$ on V&T, $+1.5\%$ on UR-FUNNY) when adding a third modality with CoMM (compared to only using two), which demonstrates its versatility. Second, we improve the state-of-the-art for SSL methods on datasets with more than two modalities ($+0.1\%$ and $+5.4\%$ on V&T and UR-FUNNY, respectively).
@@ 5 Ablation studies
Loss function. First, we check our claim that optimizing both $\mathcal{L}$ and $\sum_{i=1}^{n}\mathcal{L}_{i}$ –as in Eq. 7– is required to accurately capture uniqueness, synergy, and redundancy. In  Fig. 5, we show that minimizing $\sum_{i=1}^{n}\mathcal{L}_{i}$ improves redundancy and uniqueness, as guaranteed by our Lemma 3, but fails for synergy. Conversely, minimizing $\mathcal{L}$ allows one to learn all information terms but very slowly. In particular, for synergy, we argue this is because the model has to learn modality-specific features first (phase 1) before learning their interactions (phase 2). The former is learned through $I(Z_{i},Z^{\prime})+I(Z_{i},Z^{\prime\prime})$ while the latter is captured with $I(Z^{\prime},Z^{\prime\prime})$. Hence, $\mathcal{L}_{\text{CoMM}}$ speeds up phase 1 and can learn synergy more efficiently in phase 2.
@@
CoMM $\mathbf{99.92}_{\pm 0.03}$ $\mathbf{87.83}_{\pm 1.55}$ $71.87_{\pm 2.06}$ $\mathbf{86.83}$
@@ 6 Related work
Multimodal learning refers to methods that connect and integrate information from multiple sources of data (Baltrušaitis et al., 2018; Akkus et al., 2023). Early works focused on training separate encoders and studied different fusion mechanisms to blend the information from different inputs (Zeng et al., 2007; Pérez-Rúa et al., 2019). With the development of transformers and ViTs, the focus has shifted towards training a unique architecture and designing specific tasks and loss functions to integrate multimodal interactions (Xu et al., 2023; Lu et al., 2019; Sun et al., 2019; Chen et al., 2020b; Lu et al., 2020). Today multimodal learning includes different research lines, from generative multimodal learning (Suzuki & Matsuo, 2022; Ramesh et al., 2021; Saharia et al., 2022; Wang et al., 2022b; Alayrac et al., 2022) to multimodal representation learning (Zong et al., 2024; Tian et al., 2020a). CoMM belongs to the latter category.
@@
Self-supervised multimodal representation learning. Self-supervised learning aims to learn general representations through supervisory signals from the data itself (known as the pretext task) (Balestriero et al., 2023). In the multimodal context (Zong et al., 2024), self-supervised methods can be grouped according to their pretext task. Clustering-based methods (Alwassel et al., 2020; Asano et al., 2020), where cluster assignments are used as pseudo-labels to train the model and different modalities can be used as supervisory signals to each other; masking modeling methods (Mizrahi et al., 2023; Bachmann et al., 2022; Lu et al., 2023) that reconstruct pieces of information that have been masked from input data. In the multimodal case, masked modelling is used in a cross-modal way, by predicting missing information conditioned in other modalities. Contrastive methods (Radford et al., 2021; Jia et al., 2021) in the multimodal context have been mostly used on matched data from different modalities to obtain aligned–yet distinct–representations. In this work, instead of representing each modality separately, we take a different perspective for contrastive methods by learning a single multimodal representation of the inputs.
@@
Multimodal contrastive learning. Current methods in contrastive learning are inspired by the idea of multiview learning (Li et al., 2018) and have shown remarkable results in representation learning in unimodal tasks (Chen et al., 2020a; He et al., 2020; Tian et al., 2020a; Oord et al., 2018). The natural extension of these approaches to multimodal settings is to optimize a cross-modal contrastive loss (Radford et al., 2021; Alayrac et al., 2020; Jia et al., 2021). Other works have gone even further by introducing cross-modal and intra-modal contrastive objectives (Mu et al., 2022; Jain et al., 2021; Mi et al., 2024), or by adding other intra-modality regularization terms (Wang et al., 2022a; Kim et al., 2022). However, these approaches are designed to learn redundant information, neglecting the contributions of uniqueness or synergy. Recently, FactorCL (Liang et al., 2023b) has proposed a solution to model shared and unique task-relevant information explicitly. Yet, the method relies heavily on assumptions that are hard to meet in practice; it proposes a factorized approximation of multimodal interactions that is prone to cumulative errors and does not model synergistic information. Alternatively, we propose CoMM, a contrastive multimodal approach that leverages multimodal augmentations with a modular architecture optimized through information theory-grounded losses to capture unique, redundant and synergistic interactions.